OpenStudy (deadshot):

What is the period of the function \[f(x)=\sin \frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

What is the basic form of a sine function?

OpenStudy (deadshot):

\[\frac{ opposite }{ hypotenuse }\]

OpenStudy (jdoe0001):

y = A sin( Bx + C ) + D B = period divisor for original period of function, so if sine has an original period of \(\bf 2\pi\quad \textit{new period is }\quad \cfrac{2\pi}{B}\)

OpenStudy (deadshot):

Thanks!

OpenStudy (jdoe0001):

yw